OpenStudy (njc27):

(3p^3)^2 X 4p^7 divided by 2(p^4)^3 =

Directrix (directrix):

18p ?

OpenStudy (njc27):

thanks I got 18p^8 ??

Directrix (directrix):

Is this the question:

Directrix (directrix):

@Vengeance --> Should your numerator be 36 p ^13 and then the denominator as is. I keep getting 18 p.

OpenStudy (njc27):

sounds good to me thanks. New to this and still learning

OpenStudy (anonymous):

@Directrix-Yeah..7+6=13 omg epic fail..yeah its 18p

Directrix (directrix):

@njc27 ---> Do you want us to work out the steps of the problem here? Just say the word.

OpenStudy (njc27):

@directrix that would be great because now i am confused would you add the 7 & 6 and not multiply the sum is (3p^3)^2*4p^7/2(p^4)^3

Directrix (directrix):

I'll draw out the solution. First, a word. On a power of a power problem such as (x ^ 2 )^ 3, the exponents are multiplied to yield x^6 as the answer. This is because (x ^ 2) ^3 means the product of three x^2 terms or x^2 * x^2 * x^2 = x ^ (2+2+2) = x^6 power.

OpenStudy (anonymous):

\[(3p ^{3})^{2}*4p ^{7}/2(p ^{4})^{3}\] \[9p ^{2*3}*4p ^{7}/2p ^{4*3}\] \[9p ^{6}*4p ^{7}/2p ^{12}\] \[36p ^{6+7}/2p ^{12}\] \[36p ^{13}/2p ^{12}\] \[18p ^{13-12}\] \[18p ^{1}\] \[18p\] If anything is wrong now please kill me.

OpenStudy (anonymous):

Rules of exponents:\[a ^{m}*a ^{n}=a ^{m+n}\] \[(a ^{m})^{n}=a ^{m*n}=a ^{mn}\] \[a ^{m}/a ^{n}=a ^{m-n}\]

Directrix (directrix):

x^2 * x^3 is not a power of a power problem. The exponents are added because x^2 * x^3 means x*x* x*x*x = the product of five x terms or x^5. Note that x can be understood as x^1.

Directrix (directrix):

@ Vengeance --> It's fine.

OpenStudy (njc27):

Thank you very much that make a lot more sense now you are both wonderful. I'll try the next ones on my own. Just needed to be shown.

Directrix (directrix):

After you do one, post the question and the answer and we will check, if you like.

OpenStudy (anonymous):

Thanks Directrix, and good idea. We can double check the answers.